Can someone explain where the 1!,4!,2!,5!, etc is coming from?

Hypergeometric Probability Distribution QA=D_64)(3) _G)G). G).|5121) f(x) = W (13) .4773 12! 3!9! 220
Answer #1

Explanation

Hypergeometric distribution

The hypergeometric distribution is a discrete probability distribution that describes the probability of k successes (random draws for which the object drawn has a specified feature) in n draws, without replacement, from a finite population of size N that contains exactly K objects with that feature, wherein each draw is either a success or a failure.

f(x) which is provided in the question is its pmf - Probability mass function

In its pmf, COMBINATION IDENTITIES are used.

They can also be written as nCr. Formula for calculating combination is written in the attached file.

Factorial

Combination identities are expanded or solved using factorials which work in following way.

n! = n*(n-1)*(n-2)*........*1

6! = 6*5*4*3*2*1

Please see the attached 2 files.

You may give Thumb up if you find the answer helpful. Thanks.

In given distribution Problem, combination identities are working as fe) = (2) Nr xx WCn-x hore or NCn N 26 Since combinatioCr= (2) no r= To Calculate combinations, we use the formula n! r!(n-1)! whore n= ofitems no of items being chosen at a time.